OpenStudy (anonymous):

Hooke's law states that the force F required to stretch a spring x units beyond its natural length is directly proportional to x. (a) Express F as a function of x by means of a formula that involves a constant of proportionality k. F=kx (b) A weight of 3 pounds stretches a certain spring from its natural length of 10.2 inches to a length of 10.9 inches. Find the value of k in part (a). (c)What weight will stretch the spring in part (b) to a length of 13.7 inches?

OpenStudy (anonymous):

Shall we work on this together?

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

F=kx. In part b you are given F=3lb and the extension, x= 10.9-10.2 = 0.7in Find k from this. What do u get?

OpenStudy (anonymous):

Sorry, didn't see ya reply. So for part a, I do agree with your answer. So, for part b, we have three given values: 3 lbs 10.2 inch 10.9 inches. So the formula F = kx ... what value are we looking for in part b?

OpenStudy (anonymous):

(b) 4.29?

OpenStudy (anonymous):

Yes good job. In part c the extension, x = 13.7-10.2 = 3.5in and k=4.29. Use the formula to find F. What do you get?

OpenStudy (anonymous):

ok i just didnt kno wat to do with the two nmbrs for the length

OpenStudy (anonymous):

Extension is the length that the spring stretches from its original length (10.2in). What did u get as the Force, F ?

OpenStudy (anonymous):

15.015

OpenStudy (anonymous):

Yeppers. The great thing about physics is that *usually* if you have only one variable dealing with a specific unit of measurement (in this case length), but two or more measurements, all you have to do is subtract or add them to get one variable that you can plug in :D

OpenStudy (anonymous):

Good Job :)

OpenStudy (anonymous):

And don't forget the units. lbs